QUESTION 9 GDP Imports $440 450 50 550 600 50 50 50 640 700 All figures in the table are in billions of dollars of dollars. If exports should decrease by $20 billion at each level of GDP, other factors constant, then the equilibrium GDP for the economy will be $500 billion $650 billion 。$550 billion O $450 billion
Answer #1

Option (4).

When exports decrease by $20B,

Net exports (NX) ($B) = Exports - Imports = (50 - 20) - 60 - 30 - 60 = -30

In equilibrium, GDP = Aggregate expenditure (Closed economy) + NX = Aggregate expenditure (Closed economy) - 30

This relationship is satisfied when GDP = $450B and Aggregate expenditure (Closed economy) = $480 B.